# The SOFIA Massive (SOMA) Radio Survey. II. Radio Emission from High-Luminosity Protostars

### v1

@@ Abstract
We present centimeter continuum observations of seven high luminosity massive protostars and their surroundings as part of the SOFIA Massive (SOMA) Star Formation Survey. With data from the Very Large Array and the Australia Telescope Compact Array, we analyze the spectral index, morphology and multiplicity of the detected radio sources. The high sensitivity and high resolution observations allow us to resolve many sources. We report thirteen new detections and two previously known detections that we observed for the first time in radio frequencies. We use the observations to build radio spectral energy distributions (SEDs) to calculate spectral indices. With radio morphologies and the spectral indices, we give assessments on the nature of the sources, highlighting five sources that display a radio jet-like morphology and a spectral index consistent with ionized jets. Combining with the SOMA Radio I sample, we present the radio - bolometric luminosity relation, especially probing the regime from to . Here we find a steep rise in radio luminosity, which is expected by models that transition from shock ionization to photoionization.
@@ 1 Introduction
Massive stars impact many processes in astrophysics, yet crucial aspects of their formation remain uncertain due to observational difficulties given by their relative rarity and large distances (typically kpc), crowded environments and high extinctions (Tan et al., 2014). Various theories have been proposed to explain the formation of massive stars, but there is no general consensus on the basic nature of the formation mechanism. Theories range from the Turbulent Core Accretion (TCA) model (McKee & Tan, 2003) to Competitive Accretion (e.g., Bonnell et al., 2001; Wang et al., 2010; Grudić et al., 2022) to Protostellar Collisions (Bonnell et al., 1998).
@@
High-sensitivity centimeter continuum observations can provide valuable understanding of the earliest and most obscured phases of high-mass star formation. Tanaka et al. (2016) (hereafter TTZ16) developed models to predict the ionization structures and centimeter continuum emission properties using initial parameters from a physical model based on TCA, including infall envelope, disk and outflow properties and their predicted thermal spectral energy distribution (SED) (Zhang & Tan, 2011; Zhang et al., 2013, 2014b; Zhang & Tan, 2018). Rosero et al. (2016, 2019b) demonstrated that many centimeter continuum sources have morphologies and parameters indicative of ionized jets, which is consistent with the results from the free-free emission models developed by TTZ16. This makes radio continuum emission highly relevant for constraining the ionizing luminosity of a protostar and helping to develop and test theoretical models for the evolutionary sequence of massive star formation.
@@
In order to carry out such tests one needs a sample of massive protostars that ideally spans a range of core properties, evolutionary stages and environments. The SOFIA Massive (SOMA) Star Formation Survey (PI: J. Tan) was developed with such goals in mind. It aims to characterize a sample of 50 high and intermediate-mass protostars over a range of initial core masses, evolutionary stages and environments using SOFIA-FORCAST 10–40 m data. In Paper I of the survey (De Buizer et al., 2017) (hereafter SOMA I), the first eight sources were presented, which were mostly massive protostars. In Paper II (Liu et al., 2019) (hereafter SOMA II), seven high-luminosity sources were presented, representing some of the most massive protostars in the survey. In Paper III (Liu et al., 2020) (hereafter SOMA III), 14 intermediate-mass sources were presented. In Paper IV (Fedriani et al., 2023) (hereafter SOMA IV), 11 isolated sources, based on the 37 m imaging, were presented. In Paper V, Telkamp et al. (2025) (hereafter SOMA V) seven regions of relatively clustered massive star formation were studied.
@@
The SOMA Radio Survey was developed with the goal of obtaining cm radio continuum observations of the SOMA sample, i.e., to have more comprehensive wavelength coverage of “extended spectral energy distributions” (E-SEDs) that constrains ionized gas properties and thus help break degeneracies that arise in infrared-only SED inferred properties from the standard SOMA analysis based on the Zhang & Tan (2018) (hereafter ZT18) TCA model grid. In SOMA Radio Paper I (Rosero et al., 2019a) (hereafter SOMA Radio I), the first eight sources of SOMA I were observed with the Karl G. Jansky Very Large Array (VLA) and the centimeter continuum data presented and analyzed, including comparison to the TTZ16 models for free-free emission.
@@
Our approach follows the general methods developed in SOMA Radio I to analyze the data. In particular, we construct radio SEDs, by reducing the centimeter continuum observations, measuring the fluxes from the images, and analyzing the morphology and multiplicity of each source. We note that due to improvements in the infrared SED analysis, including a new algorithm to calculate the optimal aperture size in an unbiased and reproducible way for extended sources, updates in the SED fitting tool and revised methods assessing uncertainties in background subtracted fluxes, we use the infrared inferred results (e.g., bolometric luminosities) from SOMA V, rather than the results from SOMA II.
@@ 2 Methods
The SOMA Star Formation Survey sample is defined by SOFIA-FORCAST observations (i.e., from 7 to m). The sample we analyze in this paper are part of the protostars presented by SOMA II, as well as sources IRAS 16562-3959 N and G305.20+0.21A that were reported in SOMA III, as an addendum to the sources in the SOMA II regions. Thus a total of nine protostars in seven target regions will be analyzed: G45.12+0.13, G309.92+0.48, G35.58-0.03, IRAS 16562-3959, IRAS 16562-3959 N, G305.20+0.21, G305.20+0.21 A, G49.27-0.34 and G339.88-1.26. The radio observations presented here are our own VLA observations for regions G45.12+0.13, G35.58-0.03 and G49.27-0.34. For G305.20+0.21 and G309.92+0.48 we present our own ATCA observations, and for regions IRAS 16562-3959 and G339.88-1.26, we used ATCA observations reported by Guzmán et al. (2016) and Purser et al. (2016), respectively.
@@
Images obtained from Andrés Guzmán (via public personal archive).
Images obtained from Simon Purser (via private communication).
@@ 2.1.1 The 6 cm data
The 6 cm (C-Band) observations were carried out in the A configuration providing angular resolutions . The project code for the observations of regions G35.58-0.03 and G49.27-0.34 is 18A-294 (PI: V. Rosero) and for region G45.12+0.13 is 19A-216 (PI: V. Rosero). The data consists of two 2 GHz wide basebands (3 bit samplers) centered at 5.03 and 6.98 GHz. The data were recorded in 30 unique spectral windows (SPWs), each comprised of 64 channels and each channel being 2 MHz wide, resulting in a total bandwidth of 3842 MHz (before “flagging”). Source 3C48 was used as flux density and bandpass calibrator for regions G35.58-0.03 and G49.27-0.34, and 3C286 for region G45.12+0.13.
The data were processed using NRAO’s Common Astronomy Software Applications (CASA) package and we use calibrated data from the VLA Calibration Pipeline version 41154 (Pipeline-CASA51-P2-B) for regions G35.58-0.03 and G49.27-0.34, and the VLA Calibration Pipeline version 42270 (Pipeline-CASA54-P2-B) for region G45.12+0.13. The images were made using the tclean task and Briggs Robust = 0.5 weighting (Briggs, 1995). Imaging of regions G35.58-0.03 and G45.12+0.13 also included one round of phase-only self calibration using solution intervals ranging from 30s to 60s depending on frequency.
@@
The 1.3 cm (K-Band) observations were made in the B configuration providing angular resolutions . The project code for all observations in this band is 19A-216 (PI: V. Rosero). The data consists of two 4 GHz wide basebands (3 bit samplers) centered at 20.4 and 24.4 GHz. The data was recorded in 60 unique SPWs, comprised of 64 channels and each channel being 2 MHz wide, resulting in a total bandwidth of 7680 MHz, before “flagging”. Source 3C286 was used as flux density and bandpass calibrator. The data reduction was done in the same manner as that for the C-Band observations, using the VLA Calibration Pipeline version 42270 (Pipeline-CASA54-P2-B) for all the regions.
@@ 2.2 ATCA data
For regions G309.92+0.48, IRAS 16562-3959, G305.20+0.21 and G339.88-1.26 we used ATCA data. The project code of the observations for the regions G309.92+0.48 and G305.20+0.21 is C3396 (PI: V. Rosero). For regions IRAS 16562-3959 and G339.88-1.26, we used ATCA observations reported by Guzmán et al. (2016) and Purser et al. (2016).
The observations from project code C3396 were made using ATCA in the 6A and 6B configurations on October 2020. Observations were made at four frequencies: 5.5, 9.0, 17.0 and 22.8 GHz, each with a bandwidth of 2048 MHz (XX, YY, XY, and YX polarizations) that was split into channels of width 1 MHz.
Individual scan times on flux and phase calibrators were 5 minutes and 90 seconds respectively and on target sources were between 3 and 45 minutes dependent on atmospheric conditions and observing frequencies. The average total integration time on each science target was 75 minutes. The bandpass and flux density was calibrated with source B1934-638 using the Stevens-Reynolds 2016 model (Partridge et al., 2016). The ATCA data were calibrated in CASA following standard interferometric procedures for flagging, flux-density transfer and phase referencing. Imaging of region G305.20+0.21 also included one round of phase-only self calibration using solution intervals ranging from 30s to 60s depending on frequency.
@@
For region IRAS 16562-3959 we used the data from Guzmán et al. (2016) at 5, 9, 17 and 19 GHZ: more information about the observations can be found in section 2 of their paper. Meanwhile, for region G339.88-1.26 we used data from Purser et al. (2016) at 5.5, 9.0, 17.0 and 22.8 GHz: more information about the observations can be found in section 3.1 of their paper. In order to have frequencies comparable to the combined images made for the other sources, we select the lowest and highest frequency maps of both regions for use in the direct comparison with the combined maps.
@@ 3 Results
Following the methods of SOMA Radio I, we consider a radio source being detected when the peak intensity is at least 5 times the rms () in either of the baseband-combined images (see Section 2.1 and 2.2) at the different frequencies. For non-detections in the combined images, we report a 3 limiting value for the flux density at a given frequency.
Figure 1 shows the radio contour plots of the lower frequency in red (C-band (6 cm) for the VLA observations and 5, 5.5 or 7.2 GHz for ATCA observations) and the higher frequency in cyan (K-band (1.3 cm) for the VLA observations and 17, 19 or 19.9 GHz for ATCA observations) for all the radio sources detected in our sample and overlaid to SOFIA-FORCAST 37 m images.
@@
Tables 4-12 report the radio parameters for each of the nine protostars studied in this paper. These parameters were measured based on different size scales, as follows. The SOMA scale refers to the size of the aperture radius used in SOMA II and SOMA V to measure IR fluxes. The Intermediate scale is based on the morphology of the radio source, specifically whether the detections appear to be of jet-like nature. The Inner scale is given by the size of the central radio detection that is likely most closely associated with the driving protostar (i.e., association with compact millimeter dust continuum emission). For the VLA data presented, we determined the flux density in each wideband image in the SOMA and Intermediate scales by using the task imstat of CASA, either enclosing the SOMA aperture in a circular region or enclosing the elongated jet-like structure in a box, respectively (see Table 13). The uncertainties of the flux densities for these two scales are estimated as (npts/beam area)0.5 added in quadrature with an assumed 10% error in calibration, where is the rms of the image, npts is the number of pixels enclosed in the box or the circular region, and the beam area is the number of pixels within a synthesized beam of the image. For the Inner scale, we determined the flux density using the task imfit of CASA, and the uncertainty was estimated using the statistical error from the Gaussian fit added in quadrature with an assumed 10% error in calibration.
@@
Column 1 of Tables 4-12 shows the given scale; for each, columns 2 and 3 report the R.A. and Decl. For the SOMA scale, these refer to the pointing center observations of SOFIA-FORCAST (see SOMA II Table 1); for the Intermediate scale, they refer to a middle point in the jet-like detection; and for the Inner scale, they refer to the R.A. and Decl. of the peak intensity of the central detected object. The following columns are the flux densities () at different frequencies, with the uncertainties given in parentheses (for non-detections, the upper limits denotes the minimum detectable flux of a point source). The last column in Tables 4-12 report the spectral indices and their uncertainties at each scale (see Section 3.2). Since the radio data are not sensitive to extended emission over scales as large as the SOMA and possibly the Intermediate scales, the flux measurements represent the sum over all compact sources within the scale, and the spectral indices contain some corresponding uncertainty (more details in Section 3.2). The error bars for these measurements are large, due to having many independent beams within the scale.
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 19:13:27.86 | +10.53.36.60 | 1797.44(180.6) | 1855.93(186.5) | 3673.00(367.5) | 3593.19(360.71) | 0.5(0.1) |
Intermediate | |||||||
Inner | 19:13:27.87 | +10.53.36.17 | 1702.69(51.11) | 1928.91(57.89) | 3668.96(110.0) | 3722.08(111.7) | 0.5(0.1) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 13:50:41.85 | 61.35.10.40 | 454.76(47.4) | 753.47(79.0) | 1017.68(108.6) | 955.54(115.6) | 0.5(0.1) |
Intermediate | |||||||
Inner | 13:50:41.84 | 61.35.10.85 | 428.18(14.6) | 713.91(29.3) | 1085.90(57.86) | 1191.15(76.69) | 0.7(0.1) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 18:56:22.52 | +02.20.27.20 | 172.38(17.3) | 199.20(20.0) | 302.12(30.6) | 321.61(33.5) | 0.4(0.1) |
Intermediate | |||||||
Inner a | 18:56:22.52 | +02.20.30.53 | 84.00(27.32) | 118.00(35.05) | 237.00(91.14) | 263.00(101.47) | 0.7(0.2) |
@@
Note. — The Intermediate scale corresponds to the extent of the radio jet. Units of R.A. are hours, minutes, and seconds. Units of Decl. are degrees, arcminutes, and arcseconds.

### v3

@@ Abstract
We present centimeter continuum observations of seven high luminosity massive protostars and their surrounding sources in regions with multiple targets, as part of the SOFIA Massive (SOMA) Star Formation Survey. With data from the Very Large Array and the Australia Telescope Compact Array, we analyze the spectral index, morphology and multiplicity of the detected radio sources. The high-sensitivity, high-resolution observations allow us to resolve many sources; 65 of the reported sources are resolved at least within the synthesized beam. We report fifteen new detections, thirteen of which are entirely new and two that have counterparts at other wavelengths but are detected here for the first time at radio frequencies. We use the observations to build radio spectral energy distributions (SEDs) to calculate spectral indices. With radio morphologies and the spectral indices, we give assessments on the nature of the sources, highlighting six sources that display a radio jet-like morphology and a spectral index consistent with ionized jets. Combining with the SOMA Radio I sample, we present the radio - bolometric luminosity relation, especially probing the regime from to . Here we find a steep rise in radio luminosity, which is expected by models that transition from shock ionization to photoionization.
show]sequeiramuri@wisc.edu
show]vrosero@caltech.edu
@@ 1 Introduction
Massive stars impact many processes in astrophysics, yet crucial aspects of their formation remain uncertain due to observational difficulties given by their relative rarity and large distances (typically kpc), crowded environments and high extinctions (J. C. Tan et al., 2014). Various theories have been proposed to explain the formation of massive stars, but there is no general consensus on the basic nature of the formation mechanism. Theories range from the Turbulent Core Accretion (TCA) model (C. F. McKee & J. C. Tan, 2003) to Competitive Accretion (e.g., I. A. Bonnell et al., 2001; P. Wang et al., 2010; M. Y. Grudić et al., 2022) to Protostellar Collisions (I. A. Bonnell et al., 1998).
@@
High-sensitivity centimeter continuum observations can provide valuable understanding of the earliest and most obscured phases of high-mass star formation. K. E. I. Tanaka et al. (2016) (hereafter TTZ16) developed models to predict the ionization structures and centimeter continuum emission properties using initial parameters from a physical model based on TCA, including infall envelope, disk and outflow properties and their predicted thermal spectral energy distribution (SED) (Y. Zhang & J. C. Tan, 2011; Y. Zhang et al., 2013, 2014; Y. Zhang & J. C. Tan, 2018). V. Rosero et al. (2016, 2019b) demonstrated that many centimeter continuum sources have morphologies and parameters indicative of ionized jets, which is consistent with the results from the free-free emission models developed by TTZ16. This makes radio continuum emission highly relevant for constraining the ionizing luminosity of a protostar and helping to develop and test theoretical models for the evolutionary sequence of massive star formation.
@@
In order to carry out such tests one needs a sample of massive protostars that ideally spans a range of core properties, evolutionary stages and environments. The SOFIA Massive (SOMA) Star Formation Survey (PI: J. Tan) was developed with such goals in mind. It aims to characterize a sample of 50 high and intermediate-mass protostars over a range of initial core masses, evolutionary stages and environments using SOFIA-FORCAST 10–40 m data. In Paper I of the survey (J. M. De Buizer et al., 2017) (hereafter SOMA I), the first eight sources were presented, which were mostly massive protostars. In Paper II (M. Liu et al., 2019) (hereafter SOMA II), seven high-luminosity sources were presented, representing some of the most massive protostars in the survey. In Paper III (M. Liu et al., 2020) (hereafter SOMA III), 14 intermediate-mass sources were presented. In Paper IV (R. Fedriani et al., 2023) (hereafter SOMA IV), 11 isolated sources, based on the 37 m imaging, were presented. In Paper V, Z. Telkamp et al. (2025) (hereafter SOMA V) seven regions of relatively clustered massive star formation were studied.
@@
The SOMA Radio Survey was developed with the goal of obtaining cm radio continuum observations of the SOMA sample, i.e., to have more comprehensive wavelength coverage of “extended spectral energy distributions” (E-SEDs) that constrains ionized gas properties and thus help break degeneracies that arise in infrared-only SED inferred properties from the standard SOMA analysis based on the Y. Zhang & J. C. Tan (2018) (hereafter ZT18) TCA model grid. In SOMA Radio Paper I (V. Rosero et al., 2019a) (hereafter SOMA Radio I), the first eight sources of SOMA I were observed with the Karl G. Jansky Very Large Array (VLA) and the centimeter continuum data presented and analyzed, including comparison to the TTZ16 models for free-free emission.
@@
Our approach follows the general methods developed in SOMA Radio I to analyze the data. In particular, we construct radio SEDs by reducing the centimeter continuum observations, measuring the fluxes from the images, and analyzing the morphology and multiplicity of each region. We define multiplicity in a region when it contains two or more 5 detections that are not necessarily associated with each other. We note that due to improvements in the infrared SED analysis, including a new algorithm to calculate the optimal aperture size in an unbiased and reproducible way for extended sources, updates in the SED fitting tool and revised methods assessing uncertainties in background subtracted fluxes, we use the infrared inferred results (e.g., bolometric luminosities) from SOMA V, rather than the results from SOMA II.
@@ 2 Methods
The SOMA Star Formation Survey sample is defined by SOFIA-FORCAST observations (i.e., from 7 to m). The sample we analyze in this paper are part of the protostars presented by SOMA II, as well as sources IRAS 16562-3959 N and G305.20+0.21A that were reported in SOMA III, as an addendum to the sources in the SOMA II regions. Thus a total of nine protostars in seven target regions will be analyzed: G45.12+0.13, G309.92+0.48, G35.58-0.03, IRAS 16562-3959, IRAS 16562-3959 N, G305.20+0.21, G305.20+0.21 A, G49.27-0.34 and G339.88-1.26. The radio observations presented here are our own VLA observations for regions G45.12+0.13, G35.58-0.03 and G49.27-0.34. For G305.20+0.21 and G309.92+0.48 we present our own ATCA observations, and for regions IRAS 16562-3959 and G339.88-1.26, we used ATCA observations reported by A. E. Guzmán et al. (2016) and S. J. D. Purser et al. (2016), respectively.
@@
Images obtained from Andrés Guzmán (via public personal archive). Observations reported in A. E. Guzmán et al. (2016).
Images obtained from Simon Purser (via private communication). Observations reported in S. J. D. Purser et al. (2016).
@@ 2.1.1 The 6 cm data
The 6 cm (C-Band) observations were carried out in the A configuration providing angular resolutions and Largest Angular Scale (LAS or ) of 8.9 arcsec. The project code for the observations of regions G35.58-0.03 and G49.27-0.34 is 18A-294 (PI: V. Rosero) and for region G45.12+0.13 is 19A-216 (PI: V. Rosero). The data consists of two 2 GHz wide basebands (3 bit samplers) centered at 5.03 and 6.98 GHz. The data were recorded in 30 unique spectral windows (SPWs), each comprised of 64 channels and each channel being 2 MHz wide, resulting in a total bandwidth of 3842 MHz (before “flagging”). Source 3C48 was used as flux density and bandpass calibrator for regions G35.58-0.03 and G49.27-0.34, and 3C286 for region G45.12+0.13.
The data were processed using NRAO’s Common Astronomy Software Applications (CASA) package and we use calibrated data from the VLA Calibration Pipeline version 41154 (Pipeline-CASA51-P2-B) for regions G35.58-0.03 and G49.27-0.34, and the VLA Calibration Pipeline version 42270 (Pipeline-CASA54-P2-B) for region G45.12+0.13. The images were made using the tclean task and Briggs Robust = 0.5 weighting (D. S. Briggs, 1995). Imaging of regions G35.58-0.03 and G45.12+0.13 also included one round of phase-only self calibration using solution intervals ranging from 30s to 60s depending on frequency.
@@
The 1.3 cm (K-Band) observations were made in the B configuration providing angular resolutions and Largest Angular Scale (LAS or ) of 7.9 arcsec. The project code for all observations in this band is 19A-216 (PI: V. Rosero). The data consists of two 4 GHz wide basebands (3 bit samplers) centered at 20.4 and 24.4 GHz. The data was recorded in 60 unique SPWs, comprised of 64 channels and each channel being 2 MHz wide, resulting in a total bandwidth of 7680 MHz, before “flagging”. Source 3C286 was used as flux density and bandpass calibrator. The data reduction was done in the same manner as that for the C-Band observations, using the VLA Calibration Pipeline version 42270 (Pipeline-CASA54-P2-B) for all the regions.
@@ 2.2 ATCA data
For regions G309.92+0.48 and G305.20+0.21, we used our own ATCA observations; meanwhile, for regions IRAS 16562-3959 and G339.88-1.26, we used ATCA observations reported by A. E. Guzmán et al. (2016) and S. J. D. Purser et al. (2016).
In the following, we describe our new ATCA observations. The project code of the observations for the regions G309.92+0.48 and G305.20+0.21 is C3396 (PI: V. Rosero). These observations were made using ATCA in the 6A and 6B configurations in October 2020. Observations were made at four frequencies: 5.5, 9.0, 17.0 and 22.8 GHz, each with a bandwidth of 2048 MHz (XX, YY, XY, and YX polarizations) that was split into channels of width 1 MHz.
Individual scan times on flux and phase calibrators were 5 minutes and 90 seconds respectively and on target sources were between 3 and 45 minutes dependent on atmospheric conditions and observing frequencies. The average total integration time on each science target was 75 minutes. The bandpass and flux density was calibrated with source B1934-638 using the Stevens-Reynolds 2016 model (B. Partridge et al., 2016). The ATCA data were calibrated in CASA following standard interferometric procedures for flagging, flux-density transfer and phase referencing. Imaging of region G305.20+0.21 also included one round of phase-only self calibration using solution intervals ranging from 30s to 60s depending on frequency.
@@
For region IRAS 16562-3959 we used the data from A. E. Guzmán et al. (2016) at 5, 9, 17 and 19 GHZ: more information about the observations can be found in section 2 of their paper. Meanwhile, for region G339.88-1.26 we used data from S. J. D. Purser et al. (2016) at 5.5, 9.0, 17.0 and 22.8 GHz: more information about the observations can be found in section 3.1 of their paper. In order to have frequencies comparable to the combined images made for the other sources, we select the lowest and highest frequency maps of both regions for use in the direct comparison with the combined maps, but we used the images in the same way that they were published.
@@ 3 Results
Following the methods of SOMA Radio I, we consider a radio source being detected when the peak intensity is at least 5 times the rms () in either of the baseband-combined images (see Section 2.1 and 2.2) at the different frequencies. For non-detections in the combined images, we report an upper limit value of 3 for the flux density at the given frequency.
Figure 1 shows the radio contour plots of the lower frequency in red (C-band (6 cm) for the VLA observations and 5, 5.5 or 7.2 GHz for ATCA observations) and the higher frequency in cyan (K-band (1.3 cm) for the VLA observations and 17, 19 or 19.9 GHz for ATCA observations) for all the radio sources detected in our sample and overlaid to SOFIA-FORCAST 37 m images (from SOMA II).
@@
Table 4 reports the radio parameters for each of the nine protostars studied in this paper. In line with the methods developed in SOMA Radio I, these parameters were measured based on different size scales, as follows. The SOMA scale refers to the size of the aperture radius used in SOMA II and SOMA V to measure IR fluxes. The Intermediate scale is based on the morphology of the radio source, specifically whether the detections appear to be of jet-like nature. The Inner scale is given by the size of the central radio detection that is likely most closely associated with the driving protostar, as set by the SOFIA data and the center of the infrared SEDs analyzed in SOMA II (i.e., association with compact millimeter dust continuum emission). For the VLA data presented, we determined the flux density in each wideband image in the SOMA and Intermediate scales by using the task imstat of CASA, either enclosing the SOMA aperture in a circular region or enclosing the elongated jet-like structure in a box, respectively (see Table 5). The uncertainties of the flux densities for these two scales are estimated as (npts/beam area)0.5 added in quadrature with an assumed 10% error in calibration, where is the rms of the image, npts is the number of pixels enclosed in the box or the circular region, and the beam area is the number of pixels within a synthesized beam of the image. For the Inner scale, we determined the flux density using the task imfit of CASA, and the uncertainty was estimated using the statistical error from the Gaussian fit added in quadrature with an assumed 10% error in calibration.
@@
Column 1 and 2 of Table 4 shows the region and the given scale; for each, columns 3 and 4 report the R.A. and Decl. For the SOMA scale, these refer to the pointing center observations of SOFIA-FORCAST (see SOMA II Table 1); for the Intermediate scale, they refer to a middle point in the jet-like detection; and for the Inner scale, they refer to the R.A. and Decl. of the peak intensity of the central detected object. The following columns are the flux densities () at different frequencies, with the uncertainties given in parentheses (for non-detections, the upper limits denotes the minimum detectable flux of a point source). The last column in Table 4 reports the spectral indices and their uncertainties at each scale (see Section 3.2). Since we are limited by the Large Angular Scale (LAS), the radio data are not sensitive to extended emission over scales as large as the SOMA and possibly the Intermediate scales, the flux measurements represent the sum over all compact sources within the scale, and the spectral indices contain some corresponding uncertainty (more details in Section 3.2). The error bars for these measurements are large, due to having many independent beams within the scale. Consequently, the uncertainties in the flux density and spectral index at these scales tend to be higher for some sources. It is important to note, however, that smaller structures such as the Inner scale are less affected.
| Region | Scale | R.A | Decl. | Spectral | ||||
| (J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | ||
| SOMA | 19:13:27.86 | +10:53:36.60 | 1797.44(180.6) | 1855.93(186.5) | 3673.00(367.5) | 3593.19(360.71) | 0.5(0.1) | |
| G45.12+0.13 | Intermediate | |||||||
| Inner | 19:13:27.87 | +10:53:36.17 | 1702.69(51.11) | 1928.91(57.89) | 3668.96(110.0) | 3722.08(111.7) | 0.5(0.1) | |
| SOMA | 13:50:41.85 | 61:35:10.40 | 454.76(47.4) | 753.47(79.0) | 1017.68(108.6) | 955.54(115.6) | 0.5(0.1) | |
| G309.92+0.48a | Intermediate | |||||||
| Inner | 13:50:41.84 | 61:35:10.85 | 428.18(14.6) | 713.91(29.3) | 1085.90(57.86) | 1191.15(76.69) | 0.7(0.1) | |
| SOMA | 18:56:22.52 | +02:20:27.20 | 172.38(17.3) | 199.20(20.0) | 302.12(30.6) | 321.61(33.5) | 0.4(0.1) | |
| G35.58-0.03 | Intermediate | |||||||
| Inner | 18:56:22.52 | +02:20:30.53 | 84.00(27.32) | 118.00(35.05) | 237.00(91.14) | 263.00(101.47) | 0.7(0.2) | |
| SOMA | 16:59:41.63 | 40:03:43.60 | 18.92(1.9) | 23.50(2.4) | 34.27(3.7) | 31.56(3.5) | 0.4(0.1) | |
| IRAS 16562-3959b | Intermediate | 16:59:42.17 | 40:03:46.85 | 19.66(2.0) | 27.63(2.8) | 31.84(3.2) | 33.34(3.4) | 0.4(0.1) |
| Inner | 16:59:41.63 | 40:03:43.73 | 8.18(0.87) | 13.40(1.43) | 20.80(2.19) | 22.90(2.43) | 0.8(0.1) | |
| SOMA | 16:59:43.01 | 40:03:11.56 | 0.39 | 0.47(0.22) | 0.59(0.64) | 0.87(0.66) | 0.7 | |
| IRAS 16562-3959 Nb | Intermediate | |||||||
| Inner | 0.06 | 0.06 | 0.09 | 0.08 | 0.4 | |||
| SOMA | 13:11:13.64 | 62:34:40.70 | 1.25(2.37) | 1.05(2.14) | 0.71(2.18) | 3.22(8.43) | -0.3(3.1) | |
| G305.21+0.21a | Intermediate | |||||||
| (G305A) | Inner | 13:11:13.75 | 62:34:41.31 | 0.96(0.01) | 0.74(0.04) | 0.91(0.09) | 2.70(0.27) | 0.2(0.2) |
| SOMA | 13:11:10.49 | 62:34:38.80 | 16.57(2.11) | 12.98(2.36) | 3.42(2.63) | 29.72 | -1.0 | |
| G305.20+0.21a | Intermediate | 13:11:10.52 | 62:34:39.08 | 7.23(0.83) | 4.55(0.77) | 1.59(0.83) | 9.33 | -1.2 |
| (G305B) | Inner | 13:11:10.56 | 62:34:38.31 | 3.39(0.13) | 1.61(0.89) | 1.37(0.19) | 1.92(0.38) | -0.7(0.2) |
| SOMA | 19:23:06.61 | +14:20:12.00 | 44.30(4.54) | 52.93(5.44) | 26.11(3.34) | 40.86(4.50) | -0.3(0.1) | |
| G49.27-0.34c | Intermediate | 19:23:06.69 | +14:20:11.22 | 0.21(0.04) | 0.01(0.04) | 0.33(0.08) | 0.34(0.07) | 0.7(0.5) |
| Inner | 19:23:06.66 | +14:20:11.66 | 0.04(0.01) | 0.03(0.01) | 0.27(0.02) | 0.35(0.02) | 1.6(0.2) | |
| SOMA | 16:52:04.67 | 46:08:34.20 | 5.50(0.71) | 3.64(0.68) | 4.11(2.16) | 4.27(4.61) | -0.6(0.5) | |
| G339.88-1.26a | Intermediate | 16:52:04.62 | 46:08:36.04 | 5.31(0.57) | 6.86(0.73) | 4.58(1.04) | 4.03(2.05) | -0.1(0.2) |
| Inner | 16:52:04.94 | 46:08:31.32 | 0.86(0.09) | 1.51(0.16) | 2.44(0.26) | 3.00(0.33) | 0.9(0.1) |
@@
Note. — The Intermediate scale corresponds to the extent of the radio jet (candidate). Units of R.A. are hours, minutes, and seconds. Units of Decl. are degrees, arcminutes, and arcseconds.